Let the demand function for a product be Q = 50 − 5P. The inverse demand function of this demand function is:

1) P = 10 + 0.2Q

2) P = 50 − 0.2Q

3) P = 10 − 0.2Q

4) Q = 25 + P

Please give the # of the answer.
Answer #1

Ans: 3) P = 10 − 0.2Q

Explanation:

Q = 50 − 5P

Inverse demand function , P = ( 50 / 5 ) - (1 / 5 * Q) = 10 - 0.2Q